# A Fast Kernel-based Conditional Independence test with Application to Causal Discovery

### v1

@@ 1 Introduction
Conditional independence (CI) testing is a fundamental operation in causal discovery and structure learning. Widely used algorithms such as the PC algorithm [25] and Fast Causal Inference [24] rely on CI tests to recover the causal skeleton of a graph from observational data. The core statistical question is whether two variables $X$ and $Y$ are independent given a conditioning set $Z$, that is, whether $X\perp\!\!\!\!\perp Y\mid Z$. Despite being frequently adapted to fields like neuroscience [23], climate research [7] or economics [2], CI testing remains the computational bottleneck in constraint-based causal discovery, especially as sample size increases [1, 17, 22].
Standard CI tests suit different types of data and assumptions: Traditional tests like the Fisher-$Z$ test [8] assume linear Gaussian data, while discrete tests such as $\chi^{2}$ require categorical variables. More recent approaches avoid strong distributional assumptions by leveraging nonparametric techniques such as kernel methods. Among these, the Kernel-based Conditional Independence test (KCI) [32] has become a standard choice due to its flexibility and empirical power. KCI is based on Hilbert space embeddings and computes dependence via kernel covariance operators, making it applicable to arbitrary continuous distributions.
KCI requires operations on $n\times n$ Gram matrices and matrix inversions, resulting in $\mathcal{O}(n^{3})$ runtime per test, which makes it infeasible for large-scale applications. Recent work has sought to mitigate this cost via sample splitting [19], random Fourier features [28], neural network approximations [6], and randomization-based tests [21]. However, these approximations can degrade statistical power and require additional tuning.
Our goal is to accelerate the KCI test without sacrificing its statistical rigor and nonparametric flexibility. To this end, we propose FastKCI, a novel variant that leverages ideas from embarrassingly parallel inference in Gaussian processes [34]. FastKCI partitions data based on a generative model in the conditioning set $Z$, performs KCI tests in parallel on each partition, and aggregates the test statistics using an importance weighting scheme. This blockwise strategy enables significant computational speedups, especially in multi-core or distributed environments.
@@ 2 Related Work
Due to the growing importance of causal inference, considerable research has been devoted to discovering efficient methods to identify causal structures from observational data [29]. Traditional causal discovery methods are often categorized as constraint-based or score-based. Constraint-based methods, such as the PC algorithm [25] or FCI [26], rely on CI tests to identify the underlying causal structure. Score-based methods, such as Greedy Equivalence Search [3], evaluate causal structures based on scoring criteria. Both approaches face computational challenges: constraint-based methods due to intensive CI testing, and score-based methods due to an exponentially large search space.
@@
Existing work has improved the efficiency of the PC algorithm by reducing unnecessary CI tests [27], optimizing the overall search procedure itself—through order-independent execution [4], skipping costly orientation steps [5], sparsity-aware pruning [15], divide-and-conquer partitioning [13], and parallelism of the CI tests [17, 30, 12]. These methods, however, do not address the cubic-time bottleneck of kernel-based CI tests.
@@
Attempts in the literature to address this issue take several forms. The original KCI paper already derived an analytic $\Gamma$-approximation of the null distribution and proposed simple median-heuristic bandwidth choices to avoid costly resampling [32]. Strobl et al. [28] accelerate KCIT by replacing the full kernel matrices with an $m$-dimensional random Fourier-feature approximation. Doran et al. [6] re-express conditional independence as a single kernel two-sample problem by restricting permutations of $(X,Y)$, thereby changing the test statistic while lowering runtime. Zhang et al. [31] eliminate kernel eigen-decompositions altogether by regressing $X$ and $Y$ on $Z$ and measuring residual similarity with a lightweight kernel. Additional ideas include calibrating test statistics with locality-based permutations in the conditioning set [16], evaluating analytic kernel embeddings at a finite set of landmark points [20], and controlling small-sample bias via data splitting [19]. However, these approaches can compromise statistical power, particularly under complex nonlinear dependencies and in high-dimensional conditioning sets.
@@
Our method preserves the KCI statistic by evaluating it on Gaussian-mixture strata and aggregating with importance weights. We therefore leverage techniques from Gaussian Process regression, which often faces the identical problem of poor scalability due to cubic complexity [^1] In fact, as KCI solves GP regression problems to find the RKHS bandwidth, it is a sub-problem of CI. [34]. A common solution is to assume the underlying distribution of the covariates $Z=\{z_{1},\dots,z_{j}\}$ to be a mixture-of-experts (MoE) [14]. Local approaches, as in Gramacy and Lee [10], then aggregate over multiple partitions by MCMC. Zhang and Williamson [34] propose an importance sampling approach to MoE that efficiently aggregates over multiple partitions. To the best of our knowledge, our approach—using importance-sampled partitions of the data, performing parallel kernel tests, and aggregating via importance weighting—represents the first attempt explicitly bridging embarrassingly parallel inference with scalable kernel-based CI testing.
@@ 3 Background
The KCI builds on a notion of conditional independence introduced by Fukumizu et al. [9]. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and $X\in\mathcal{X}$, $Y\in\mathcal{Y}$, $Z\in\mathcal{Z}$ random variables. For each domain we fix a measurable, bounded and characteristic kernel, e.g. the Gaussian RBF, denoted $k_{X},k_{Y},k_{Z}$, with corresponding reproducing-kernel Hilbert spaces (RKHS) $\mathcal{H}_{X},\mathcal{H}_{Y},\mathcal{H}_{Z}$. Feature maps are denoted $\varphi_{X}(x)=k_{X}(x,\cdot)$ etc. Expectations are shorthand $\mathbb{E}[\cdot]$, tensor products $\otimes$, and centered features $\tilde{\varphi}_{X}\!:=\!\varphi_{X}-\mu_{X}$ where $\mu_{X}=\mathbb{E}[\varphi_{X}(X)]$.
@@ 3.2 Finite-sample KCI statistic
Given $n$ observations $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$, assemble Gram matrices $K_{X},K_{Y},K_{Z}\!\in\!\mathbb{R}^{n\times n}$ with $(K_{X})_{ij}=k_{X}(x_{i},x_{j})$, $K_{Y},K_{Z}$ analogously. Let $H:=I_{n}-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\!\top}$ and define the projection onto $Z$–residuals
@@
Intuitively, $R_{Z}$ acts like a regression operator that projects any function of the samples onto the subspace of functions depending on $Z$. Thus, $I-R_{Z}$ removes components explained by $Z$. Residualized, centered kernels are $\tilde{K}_{X}=R_{Z}HK_{X}HR_{Z}$ and $\tilde{K}_{Y}$ analogously.
The KCI test statistic of Zhang et al. [32] is then given by the Hilbert–Schmidt norm of the cross-covariance between residuals:
@@
Zhang et al. [32] show that under $H_{0}$ this statistic converges in distribution to a weighted sum of $\chi^{2}$ variables. In practice one uses a finite-sample null distribution to assess significance. We follow Proposition 5 in [32], using a spectral approach to simulate the null: We compute the eigenvalues ${\lambda_{m}}$ of a normalized covariance operator associated with $\tilde{K}_{X}$ and $\tilde{K}_{Y}$, then generate null samples and set
@@
with i.i.d. $\chi_{1}^{2}$ variates $\chi_{1,m,b}^{2}$. This approximate distribution of $T_{\text{KCI}}$ under $H_{0}$ obtains a $p$-value. The full KCI procedure is given in 2.
@@ 4 FastKCI: A Scalable and Parallel Kernel-based Conditional Independence Test
To overcome the $\mathcal{O}(n^{3})$ bottleneck, we propose FastKCI, which leverages a mixture-of-experts model and importance sampling. The core idea is to break the full kernel computation into $V$ smaller pieces (experts) corresponding to data partitions, compute local CI statistics on each piece, and then recombine them to recover the global statistic. By doing so, FastKCI achieves significant speed-ups – roughly on the order of $1/V^{2}$ of the cost of KCI – while provably maintaining the test’s correctness under mild assumptions.
@@
Thus, ${z_{i}}$ are i.i.d. draws from a $V$-component Gaussian mixture with unknown means $\mu_{v}^{Z}$ and covariances $\Sigma_{v}^{Z}$. We place a weak Normal-Inverse-Wishart prior on these parameters to allow uncertainty. This mixture-of-experts prior on $Z$ guides partitioning of the dataset. Intuitively, if $Z$ has a multimodal or complex distribution, this model lets us divide the data into $V$ clusters $\mathcal{C}_{1},\dots,\mathcal{C}_{V}$ (with $\mathcal{C}_{v}=\{i\mid U_{i}=v\}$) such that within each $\mathcal{C}_{v}$, $Z$ is roughly Gaussian. We do not assume anything restrictive about $X$ and $Y$ globally. By conditioning on $\mathcal{C}_{v}$, the relationship between $X$ and $Y$ can be analyzed locally. In particular, conditioned on a given partition $U_{1:n}=(U_{1},\dots,U_{n})$, the kernel matrices $K_{X},K_{Y},K_{Z}$ acquire an approximate block structure: after permuting indices by cluster, each matrix breaks into $V$ blocks (sub-matrices) corresponding to points in the same cluster, with negligible entries for cross-cluster pairs, especially if clusters are well-separated in $Z$. Each cluster $v$ defines a local sub-problem of size $n_{v}=|\mathcal{C}_{v}|$, and within that cluster we can perform a CI test on the restricted data $\{(x_{i},y_{i},z_{i}):i\in\mathcal{C}_{v}\}$. This yields a local test statistic $T_{v}$ and local null distribution for cluster $v$. By appropriately combining these local results, we can recover a valid global test statistic without ever computing the full $n\times n$ kernel on all data at once. In the next paragraphs, we explain the procedure in detail.
Figure 1: Motivation of the partitioning scheme in the data. The full covariance kernel estimation (left) is inefficient, while creating a partition (middle) may neglect some covariance structure. We propose to use multiple partitions (right) in parallel across $J$ parallel cores. We combine them using importance sampling. The figure is inspired by Zhang [33].
@@
As visible in Figure 1, for $j=1,\dots,J$ we independently draw a cluster assignment $U^{(j)}_{1:n}=(U_{1}^{(j)},\dots,U_{n}^{(j)})$ according to the MoE model on $Z$. The mixture prior is fit to the empirical $Z$ data. For example, one iteration $j$ samples mixture parameters $(\mu_{1}^{Z},\dots,\mu_{V}^{Z},\Sigma_{1}^{Z},\dots,\Sigma_{V}^{Z})\sim P(\mu,\Sigma)$ from the NIW prior and mixture weights $\pi$, then assigns each point $i$ to a cluster $U_{i}^{(j)}=v$ with probability
@@
For each partition $j$ and for each cluster $v\in\{1,\dots,V\}$, let $\mathcal{C}_{v}^{(j)}=\{i\mid U_{i}^{(j)}=v\}$ be indices in the partition of size $n_{v}^{(j)}=|\mathcal{C}_{v}^{(j)}|$. We form $n_{v}^{(j)}\times n_{v}^{(j)}$ Gram matrices $K_{X}^{(j,v)}$, $K_{Y}^{(j,v)}$, and $K_{Z}^{(j,v)}$. $\tilde{K}_{X}^{(j,v)}$ and $\tilde{K}_{Y}^{(j,v)}$ are calculated analogously to Equation 1 with local regression operators. The local test statistic follows as
@@
and the combined null distribution is taken as the mixture of all partition null samples with the same weights. In practice, we merge the $J$ sets of null samples $\{T_{\text{null}}^{(j,b)}\}$ into one weighted empirical distribution. Specifically, we compute the weighted empirical cumulative density $F_{T,\text{null}}(t)=\sum_{j=1}^{J}w_{j}\Big{(}\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}\{T_{\text{null}}^{(j,b)}\leq t\}\Big{)}$, which is a mixture of the $J$ null distributions.

### v2

@@ 1 Introduction
Conditional independence (CI) testing is a fundamental operation in causal discovery and structure learning. Widely used algorithms such as the PC algorithm (spirtes1991) and Fast Causal Inference (spirtes2001) rely on CI tests to recover the causal skeleton of a graph from observational data. The core statistical question is whether two variables $X$ and $Y$ are independent given a conditioning set $Z$, that is, whether $X\perp\!\!\!\!\perp Y\mid Z$. Despite being frequently adapted to fields like neuroscience (smith2011), climate research (ebertuphoff2012) or economics (awokuse2003), CI testing remains the computational bottleneck in constraint-based causal discovery, especially as sample size increases (agarwal2023; Le2019; shiragur2024).
Standard CI tests suit different types of data and assumptions: Traditional tests like the Fisher-$Z$ test (fisher1921) assume linear Gaussian data, while discrete tests such as $\chi^{2}$ require categorical variables. More recent approaches avoid strong distributional assumptions by leveraging nonparametric techniques such as kernel methods. Among these, the Kernel-based Conditional Independence test (KCI) (zhang2012) has become a standard choice due to its flexibility and empirical power. KCI is based on Hilbert space embeddings and computes dependence via kernel covariance operators, making it applicable to arbitrary continuous distributions.
KCI requires operations on $n\times n$ Gram matrices and matrix inversions, resulting in $\mathcal{O}(n^{3})$ runtime per test, which makes it infeasible for large-scale applications. Recent work has sought to mitigate this cost via sample splitting (pogodin2024), random Fourier features (Strobl2018), neural network approximations (doran2014), and randomization-based tests (shah2020). However, these approximations can degrade statistical power and require additional tuning.
Our goal is to accelerate the KCI test without sacrificing its statistical rigor and nonparametric flexibility. To this end, we propose FastKCI, a novel variant that leverages ideas from embarrassingly parallel inference in Gaussian processes (zhang2020). FastKCI partitions data based on a generative model in the conditioning set $Z$, performs KCI tests in parallel on each partition, and aggregates the test statistics using an importance weighting scheme. This blockwise strategy enables significant computational speedups, especially in multi-core or distributed environments.
@@ 2 Related Work
Due to the growing importance of causal inference, considerable research has been devoted to discovering efficient methods to identify causal structures from observational data (Zanga2022). Traditional causal discovery methods are often categorized as constraint-based or score-based. Constraint-based methods, such as the PC algorithm (spirtes1991) or FCI (spirites1995), rely on CI tests to identify the underlying causal structure. Score-based methods, such as Greedy Equivalence Search (Chickering2002), evaluate causal structures based on scoring criteria. Both approaches face computational challenges: constraint-based methods due to intensive CI testing, and score-based methods due to an exponentially large search space.
@@
Existing work has improved the efficiency of the PC algorithm by reducing unnecessary CI tests (steck1999bayesian), optimizing the overall search procedure itself—through order-independent execution (colombo2014), skipping costly orientation steps (colombo2012), sparsity-aware pruning (kalisch2007), divide-and-conquer partitioning (huang2022), and parallelism of the CI tests (Le2019; zarebavani2020; hagedorn2022). These methods, however, do not address the cubic-time bottleneck of kernel-based CI tests.
@@
Attempts in the literature to address this issue take several forms. The original KCI paper already derived an analytic $\Gamma$-approximation of the null distribution and proposed simple median-heuristic bandwidth choices to avoid costly resampling (zhang2012). Strobl2018 accelerate KCIT by replacing the full kernel matrices with an $m$-dimensional random Fourier-feature approximation. doran2014 re-express conditional independence as a single kernel two-sample problem by restricting permutations of $(X,Y)$, thereby changing the test statistic while lowering runtime. zhang2022SCIT eliminate kernel eigen-decompositions altogether by regressing $X$ and $Y$ on $Z$ and measuring residual similarity with a lightweight kernel. Additional ideas include calibrating test statistics with locality-based permutations in the conditioning set (kim2022), evaluating analytic kernel embeddings at a finite set of landmark points (scetbon2021), and controlling small-sample bias via data splitting (pogodin2024). However, these approaches can compromise statistical power, particularly under complex nonlinear dependencies and in high-dimensional conditioning sets. Concurrent with our work, guan2025 proposed an Ensemble Conditional Independence Test, which independently adopts a similar divide-and-aggregate strategy by partitioning the data, applying a generic base conditional independence test to each subset, and combining the resulting p-values via stable-distribution-based aggregation. While conceptually related, their approach and ours differ substantially in both the sample partitioning scheme and the aggregation mechanism.
@@
Our method preserves the KCI statistic by evaluating it on Gaussian-mixture strata and aggregating with importance weights. We therefore leverage techniques from Gaussian Process regression, which often faces the identical problem of poor scalability due to cubic complexity [^1] In fact, as KCI solves GP regression problems to find the RKHS bandwidth, it is a sub-problem of CI. (zhang2020). A common solution is to assume the underlying distribution of the covariates $Z=\{z_{1},\dots,z_{j}\}$ to be a mixture-of-experts (MoE) (jordan1994hierarchical). Local approaches, as in gramacy2008, then aggregate over multiple partitions by MCMC. zhang2020 propose an importance sampling approach to MoE that efficiently aggregates over multiple partitions. To the best of our knowledge, our approach – using importance-sampled partitions of the data, performing parallel kernel tests, and aggregating via importance weighting – represents the first attempt explicitly bridging embarrassingly parallel inference with scalable kernel-based CI testing.
@@ 3 Background
The KCI builds on a notion of conditional independence introduced by Fukumizu2007. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and $X\in\mathcal{X}$, $Y\in\mathcal{Y}$, $Z\in\mathcal{Z}$ random variables. For each domain we fix a measurable, bounded and characteristic kernel, e.g. the Gaussian RBF, denoted $k_{X},k_{Y},k_{Z}$, with corresponding reproducing-kernel Hilbert spaces (RKHS) $\mathcal{H}_{X},\mathcal{H}_{Y},\mathcal{H}_{Z}$. Feature maps are denoted $\varphi_{X}(x)=k_{X}(x,\cdot)$ etc. Expectations are shorthand $\mathbb{E}[\cdot]$, tensor products $\otimes$, and centered features $\tilde{\varphi}_{X}\!:=\!\varphi_{X}-\mu_{X}$ where $\mu_{X}=\mathbb{E}[\varphi_{X}(X)]$.
@@ 3.2 Finite-sample KCI statistic
Given $n$ observations $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$, assemble Gram matrices $K_{X},K_{Y},K_{Z}\!\in\!\mathbb{R}^{n\times n}$ with $(K_{X})_{ij}=k_{X}(x_{i},x_{j})$, $K_{Y},K_{Z}$ analogously. Let $H:=I_{n}-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\!\top}$ and define the projection onto ZZ–residuals
@@
Intuitively, $R_{Z}$ acts like a residual operator that removes components explained by Z, $R_{z}f$ is approximately the part of $f$ orthogonal to functions of $Z$. Residualized, centered kernels are $\tilde{K}_{X}=R_{Z}HK_{X}HR_{Z}$ and $\tilde{K}_{Y}$ analogously.
The KCI test statistic of zhang2012 is then given by the Hilbert–Schmidt norm of the cross-covariance between residuals:
@@
zhang2012 show that under $H_{0}$ this statistic converges in distribution to a weighted sum of $\chi^{2}$ variables. In practice one uses a finite-sample null distribution to assess significance. We follow Proposition 5 in zhang2012, using a spectral approach to simulate the null: We compute the eigenvalues $\lambda_{m}$} of a normalized covariance operator associated with $\tilde{K}_{X}$ and $\tilde{K}_{Y}$, then generate null samples and set
@@
with i.i.d. $\chi_{1}^{2}$ variates $\chi_{1,m,b}^{2}$. This approximate distribution of $T_{\text{KCI}}$ under $H_{0}$ obtains a $p$-value. The full KCI procedure is given in Algorithm 2.
@@ 4 FastKCI: A Scalable and Parallel Kernel-based Conditional Independence Test
To overcome the $\mathcal{O}(n^{3})$ bottleneck, we propose FastKCI, which leverages a mixture-of-experts model and importance sampling. The core idea is to break the full kernel computation into $V$ smaller pieces (experts) corresponding to data partitions, compute local CI statistics on each piece, and then recombine them to recover the global statistic. By doing so, FastKCI achieves significant speed-ups – roughly on the order of $1/V^{2}$ of the cost of KCI – while maintaining the test’s correctness under mild assumptions.
@@
Thus, $z_{i}$} are i.i.d. draws from a $V$-component Gaussian mixture with unknown means $\mu_{v}^{Z}$ and covariances $\Sigma_{v}^{Z}$. We place a weak Normal-Inverse-Wishart prior on these parameters to allow uncertainty. This mixture-of-experts prior on $Z$ guides partitioning of the dataset. Intuitively, if $Z$ has a multimodal or complex distribution, this model lets us divide the data into $V$ components $\mathcal{C}_{1},\dots,\mathcal{C}_{V}$ (with $\mathcal{C}_{v}=\{i\mid U_{i}=v\}$) such that within each $\mathcal{C}_{v}$, $Z$ is roughly Gaussian. We do not assume anything restrictive about $X$ and $Y$ globally. By conditioning on $\mathcal{C}_{v}$, the relationship between $X$ and $Y$ can be analyzed locally. In particular, conditioned on a given partition $U_{1:n}=(U_{1},\dots,U_{n})$, the kernel matrices $K_{X},K_{Y},K_{Z}$ acquire an approximate block structure: after permuting indices by cluster, each matrix breaks into $V$ blocks (sub-matrices) corresponding to points in the same cluster, with negligible entries for cross-cluster pairs, especially if clusters are well-separated in $Z$. Each component $v$ defines a local sub-problem of size $n_{v}=|\mathcal{C}_{v}|$, and within that component we can perform a CI test on the restricted data $\{(x_{i},y_{i},z_{i}):i\in\mathcal{C}_{v}\}$. This yields a local test statistic $T_{v}$ and local null distribution for component $v$. By appropriately combining these local results, we can recover a valid global test statistic without ever computing the full $n\times n$ kernel on all data at once.
The choice of a Gaussian Mixture Model assumption is motivated by its power as a universal approximator of densities, capable of modeling a wide variety of complex, multi-modal, and non-Gaussian distributions with high fidelity (McLachlan2000). The goal of FastKCI is not to assume the data is strictly Gaussian, but to leverage a flexible framework to create sensible, localized partitions of the conditioning set. Importantly, our empirical results in Section 5 (and Appendix A.1) demonstrate that FastKCI exhibits robust performance even when this assumption is misspecified or the true underlying distribution is unknown.
In the next paragraphs, we explain the procedure in detail.
Figure 1: Motivation of the partitioning scheme in the data. The full covariance kernel estimation (left) is inefficient, while partitioning the data into components a single time (middle) may neglect some of the covariance structure. We propose to use multiple partitions $J$ (right) in parallel. We combine them using importance sampling. The figure is inspired by zhang2020presentation.
@@
As visible in Figure 1, for partition sampling rounds $j=1,\dots,J$ we independently draw a cluster assignment $U^{(j)}_{1:n}=(U_{1}^{(j)},\dots,U_{n}^{(j)})$ according to the MoE model on $Z$. The mixture prior is fit to the empirical $Z$ data. In practice, we use a lightweight empirical-Bayes partitioning scheme (see Appendix B) that samples mixture components around the empirical mean of $Z$ instead of performing full EM or full posterior inference. For example, one partition sample $j$ samples mixture parameters $(\mu_{1}^{Z},\dots,\mu_{V}^{Z},\Sigma_{1}^{Z},\dots,\Sigma_{V}^{Z})\sim P(\mu,\Sigma)$ from the NIW prior and mixture weights $\pi$, then assigns each point $i$ to a component $U_{i}^{(j)}=v$ with probability
@@
For each partition $j$ and for each cluster $v\in\{1,\dots,V\}$, let $\mathcal{C}_{v}^{(j)}=\{i\mid U_{i}^{(j)}=v\}$ be indices in the cluster of size $n_{v}^{(j)}=|\mathcal{C}_{v}^{(j)}|$. We form $n_{v}^{(j)}\times n_{v}^{(j)}$ Gram matrices $K_{X}^{(j,v)}$, $K_{Y}^{(j,v)}$, and $K_{Z}^{(j,v)}$. $\tilde{K}_{X}^{(j,v)}$ and $\tilde{K}_{Y}^{(j,v)}$ are calculated analogously to Equation 1 with local regression operators. The local test statistic follows as
@@
and the combined null distribution is taken as the mixture of all partition null samples with the same weights. In practice, we merge the $J$ sets of null samples $\{T_{\text{null}}^{(j,b)}\}$ into one weighted empirical distribution. Specifically, we compute the weighted empirical cumulative density $F_{T,\text{null}}(t)=\sum_{j=1}^{J}w_{j}\Big(\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}\{T_{\text{null}}^{(j,b)}\leq t\}\Big)$, which is a mixture of the $J$ null distributions. See also Appendix B for implementation details.